Probability function of the random variable X is:

f(x)=k/2x , x = 0,1,2.......

a. Compute k.

b. Compute P(X ≥ 6).

c. Compute P(1 ≤ X ≤ 6).


1
Expert's answer
2021-09-27T14:33:52-0400

a.


"\displaystyle\sum_{x=0}^{\infin}\dfrac{k}{2^x}=k(\dfrac{1}{1-\dfrac{1}{2}})=2k=1=>k=\dfrac{1}{2}"

b.


"P(X\geq6)=1-P(X<6)"

"=1-(\dfrac{1}{2}+\dfrac{1}{4}+\dfrac{1}{8}+\dfrac{1}{16}+\dfrac{1}{32}+\dfrac{1}{64})"

"=1-\dfrac{63}{64}=\dfrac{1}{64}"

c.


"P(1\leq X\leq6)=\dfrac{1}{4}+\dfrac{1}{8}+\dfrac{1}{16}+\dfrac{1}{32}+\dfrac{1}{64}+\dfrac{1}{128}"

"=\dfrac{63}{128}"